OpenStudy (anonymous):

Whats the inverse of f(x)? f(x)=-4x^2+1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

solve \[x=-4y^2+1\] in 3 steps

OpenStudy (misty1212):

subtract 1, \[x-1=-4y^2\]

OpenStudy (misty1212):

divide by \(-4\) \[\frac{1-x}{4}=y^2\]

OpenStudy (misty1212):

take the square root don't forget the \(\pm\)

OpenStudy (nincompoop):

hmmm what happens when the function is different? how about give the concept about inverse functions :)

OpenStudy (misty1212):

what happens when the function is different?

OpenStudy (misty1212):

i guess it is a different function

OpenStudy (anonymous):

the inverse is not a function, because your function is not one to one but i guess you can still write it if you like

OpenStudy (anonymous):

still a little lost.. but i got the whole thing about switching the x and the y...

OpenStudy (anonymous):

once you switch them you need to solve for \(y\)

OpenStudy (anonymous):

\[x=-4y^2+1\]is what you get when you switch them

OpenStudy (anonymous):

then to get \(y\) you need to get rid of the 1 on the right, subtract one gives you \[x-1=-4y^2\]

OpenStudy (anonymous):

so the inverse would be.. f^-1(x)= +- sqr1-x/2

OpenStudy (anonymous):

yes it would be \[\frac{\pm\sqrt{1-x}}{2}\]

OpenStudy (anonymous):

it is not really a function, so it is a bit silly to write \[f^{-1}(x)=\frac{\pm\sqrt{1-x}}{2}\]

OpenStudy (anonymous):

btw what happened to your instagram?

OpenStudy (anonymous):

@satellite73 whatcha mean?